# White Men Lead, Black Women Help? Benchmarking Language Agency Social Biases in LLMs

### v4

@@ Abstract
Social biases can manifest in language agency. While several studies approached agency-related bias in human-written language, very limited research has investigated such biases in Large Language Model (LLM)-generated content. In addition, previous works often rely on string-matching techniques to identify agentic and communal words within texts, which fall short of accurately classifying language agency. We introduce the novel Language Agency Bias Evaluation (LABE) benchmark, which comprehensively evaluates biases in LLMs by analyzing agency levels attributed to different demographic groups in model generations. LABE leverages 5,400 template-based prompts, an accurate agency classifier, and corresponding bias metrics to test for gender, racial, and intersectional language agency biases in LLMs on 3 text generation tasks: biographies, professor reviews, and reference letters. We also contribute the Language Agency Classification (LAC) dataset, consisting of 3,724 agentic and communal sentences. Using LABE, we unveil language agency social biases in 3 recent LLMs: ChatGPT, Llama3, and Mistral. We observe that: (1) LLM generations tend to demonstrate greater gender bias than human-written texts; (2) Models demonstrate remarkably higher levels of intersectional bias than the other bias aspects. Those who are at the intersection of gender and racial minority groups—such as Black females—are consistently described by texts with lower levels of agency, aligning with real-world social inequalities; (3) Among the 3 LLMs investigated, Llama3 demonstrates the greatest overall bias; (4) Not only does prompt-based mitigation fail to resolve language agency bias in LLMs, but it frequently leads to the exacerbation of biases in generated texts.
White Men Lead, Black Women Help? Benchmarking Language Agency Social Biases in LLMs
Yixin Wan  and Kai-Wei Chang Department of Computer Science University of California, Los Angeles {elaine1wan, kwchang}@cs.ucla.edu
@@ 1 Introduction
Social biases manifest through the varying levels of agency in texts describing different demographic groups Grimm et al. (2020); Polanco-Santana et al. (2021); Stahl et al. (2022); Wan et al. (2023). For instance, bias exists in texts portraying demographic minority groups—such as Black individuals and women—as being communal (e.g. “warm” and “helpful”), and dominant social groups—such as White individuals and men—as being agentic (e.g. “authoritative” and“in charge of” things) Cugno (2020); Grimm et al. (2020). While a body of works in social science (Akos and Kretchmar, 2016; Grimm et al., 2020; Polanco-Santana et al., 2021; Park et al., 2021) and NLP (Sap et al., 2017; Ma et al., 2020; Park et al., 2021; Stahl et al., 2022; Wan et al., 2023) studied agency level in texts, they suffer from several remarkable drawbacks:
1. Existing works fail to establish a comprehensive evaluation benchmark for language agency biases in LLMs. Most aforementioned studies studied such biases in single types of human-written texts, and only focused on single dimensions of bias (e.g. only gender bias), limiting the scope of their analysis. As people are exploring more real-world downstream applications of LLM-generated texts, it is critical to identify and quantify potential agency-related fairness issues in LLM generations.
@@
2. Existing methods to measure language agency struggle with achieving accuracy and reliability. Prior works often utilized string matching for words in agentic and communal lexicons to measure agency. However, string matching and sentiment-based approaches only yield $46.65$ and $52.28$ in agency classification accuracy, respectively (as shown in Appendix B, Table 11). A qualitative example is provided in Figure 1: while differences in language agency are clearly observable in the texts, string matching yields 00 agentic and 00 communal words; sentiment classifier labels both texts as “positive”.  Wan et al. (2023) utilized a model-based method for measuring agency, but their model only achieves $66.49\%$ classification accuracy (Appendix B, Table 11).
To address these research gaps, our study proposes a novel Language Agency Bias Evaluation (LABE) benchmark for comprehensively measuring gender, racial, and intersectional language agency biases in LLMs. Using 5,400 template-based entries, an accurate language agency classifier, and interpretable metrics for each bias dimension, LABE examines agency-related biases on 3 text generation tasks for LLMs: biography, professor review, and reference letter generation. For building the accurate and reliable automated agency classification tool, we also collect and contribute the Language Agency Classification (LAC) dataset. Using a scalable, consistent, natural, and fair data collection pipeline with LLM-empowered data generation and human cross-verification, we constructed LAC with 3,724 agentic and communal sentences. Finally, we trained an agency classifier with LAC (achieving 91.69% test accuracy) and incorporated it into LABE to evaluate language agency biases in 3 recent LLMs: ChatGPT, Mistral, and Llama3. We observed that:
@@
  • Llama3 appears to be most biased among the 3 LLMs investigated. Llama3 possesses the greatest overall level of language agency bias; most severe biases in the model are observed in professor review and reference letter generation.
@@
Our LABE benchmark and LAC dataset introduce novel and valuable technical contributions, and our findings reveal previously unexplored fairness risks of LLMs from the perspective of language agency. Furthermore, we unveil the shocking fact that widely adopted prompt-based mitigation methods may intensify language agency biases in LLMs; more effective methodologies need to be developed to address the complex fairness challenge.
@@ 2.1 Biases in Human-Written and LLM-generated Texts
Bias in Reference Letters    Trix and Psenka (2003); Cugno (2020); Madera et al. (2009); Khan et al. (2021); Liu et al. (2009); Madera et al. (2019), and Wan et al. (2023) uncovered gender biases in letters of recommendation. For instance,  Trix and Psenka (2003); Madera et al. (2009) and  Madera et al. (2019) studied bias in the “exellency” of language.  Morgan et al. (2013); Akos and Kretchmar (2016); Grimm et al. (2020); Powers et al. (2020); Polanco-Santana et al. (2021); Chapman et al. (2022); Girgis et al. (2023) investigated racial biases in reference letters:  Girgis et al. (2023) studied biases in emotional words and language traits like tone, but did not open-source their evaluation tools;  Akos and Kretchmar (2016); Grimm et al. (2020); Powers et al. (2020); Chapman et al. (2022); Polanco-Santana et al. (2021), and Chapman et al. (2022) used string matching for word-level bias analysis. For example,  Powers et al. (2020) and  Chapman et al. (2022) showed that racial minority groups are significantly less frequently described with standout words than their White colleagues.
@@ 2.2 Bias in Language Agency
An increasing body of recent studies have investigated biases in intricate language styles, such as language agency Sap et al. (2017); Ma et al. (2020); Stahl et al. (2022); Wan et al. (2023).  Akos and Kretchmar (2016); Sap et al. (2017); Ma et al. (2020); Grimm et al. (2020); Polanco-Santana et al. (2021); Park et al. (2021), and Stahl et al. (2022) measured language agency by string matching for agentic and communal verbs, and then calculate their occurrence frequencies. However, string-matching methods fail to consider the diversity and complexity of language, and could not capture implicit indicators of language agency, as illustrated in Figure 1.  Wan et al. (2023) was the first to adopt a model-based method to measure language agency gender biases in LLM-generated reference letters. Nevertheless, their model lacks accuracy in sentence-level classification, and the scope of their analysis is constrained to LLM-synthesized reference letters.
@@ 3 The Language Agency Bias Evaluation (LABE) Framework
Agentic language depicts “proactive” characteristics such as speaking assertively, influencing others, and initiating tasks; communal language portrays “reactive” characteristics like caring for others, providing assistance, and sustaining relationships (Madera et al., 2009; Wan et al., 2023). We define “language agency bias” to be the unequal representation of language agency in texts depicting different demographic groups, e.g. by showing women as submissive and powerless and men as assertive and dominant (Stahl et al., 2022), or by describing racial minority groups with more communal language than agentic (Grimm et al., 2020; Polanco-Santana et al., 2021).
@@ 3.1 Generative Discriptive Texts for Demographic Groups with LLMs
Wan et al. (2023) proposed the Context-Less Generation (CLG) setting, in which they adopt templates and descriptors to prompt for a variety of LLM-generated reference letters for different genders. Inspired by CLG, we extend the setting to 3 different text generation tasks: biography, professor review, and reference letter generation. We combine descriptors with demographic information—such as race, gender, or intersectional identities—and template-based prompts to query for LLMs’ generation. Each prompt must contain race and gender descriptors. For the name descriptor, we prompt ChatGPT to generate 5 popular names for each gender and race intersectional group. Additional descriptors like occupation and department are included to improve prompt variability. The final LABE benchmark tests LLMs on 2,400 templated-based prompts for biography generation, 600 for professor review, and 2,400 for reference letters. Note that entry numbers differ due to the difference in descriptors used (departments for professor review, whereas occupations for the other 2). Full details are in Appendix A.
@@ 3.2.1 Dataset Collection
To ensure the trustworthiness of the constructed dataset, we identified and followed 4 core pillars for the data collection process:
  1. 1.
    Scalablity. Since we are constructing a classification dataset, we need to ensure enough entries for agentic language and communal language to train a useful classifier.
  2. 2.
    Consistency. Quality of texts for all entries should remain consistent, and should not differ from part to part.
  3. 3.
    Naturalness. We need to ensure that data labels align with human perceptions of “agentic” and “communal” language.
  4. 4.
    Fairness. Balancing measures should be taken to prevent potential biases from label imbalance or bias propagation from a seed dataset.
To collect a dataset with agentic and communal sentences through a mechanism that is scalable, consistent, natural, and fair, we adopt a novel dataset construction framework that consists of an automated component and a human-involved component. We begin by preprocessing a personal biography dataset (Lebret et al., 2016) into sentences, aiming at using these as seed texts to construct agentic and communal texts through paraphrasing. This step ensures the fairness of collected dataset, since (1) the raw data output would be balanced between the two labels, and (2) each sentence in each biography would have an agentic paraphrase and a communal paraphrase, preventing social bias propagation like having more agentic sentences for dominant social groups. Next, we adopt Openai’s gpt-3.5-turbo-1106 model (OpenAI, 2022) to paraphrase each sentence into an agentic version and a communal version. This ensures scalability through an automated generation pipeline, and also guarantees consistency since all paraphrases would come from a single source (in contrast with using human-written paraphrases, which is hard to scale and might result in drastically subjective writing tones). Furthermore, we utilize a human verification step to ensure the naturalness of the generated dataset. We invite $2$ human annotators, who are native speakers of English, to re-label each data and identify ambiguous cases. Finally, data entries with ambiguity are removed and ground truth labels of the LAC dataset are decided by a majority vote between the annotators’ labels and the paraphrasing target (i.e. whether a sentence was generated as an “agentic” or “communal” paraphrase). Full details of dataset construction are in Appendix B. Details on dataset statistics are in Appendix B.5
@@ 4 Unveiling Language Agency Biases in LLMs with LABE
We utilize LABE to conduct experiments on measuring gender, racial, and intersectional biases in 3 recent LLMs: ChatGPT (OpenAI, 2022), Llama3 (Touvron et al., 2023), and Mistral (AI, 2023).
Models and Generation Settings   We experiment with 3 recent LLMs: the gpt-3.5-turbo-1106 version of OpenAI s ChatGPT (OpenAI, 2022), Llama3-8B-Instruct (Touvron et al., 2023), and Mistral-7B-Instruct-v0.2. We utilize ChatGPT’s API for experiments, with no license information. Llama3 is licensed under the Meta Llama 3 Community License and Mistral is under Apache License 2.0; both models are publicly available. For ChatGPT, we followed all default generation settings in the API call. We use Huggingface’s text generation pipeline to implement Llama3 and Mistral, and follow all default generation hyperparameters besides setting maximum number of new tokens to 512. All results are averaged on random seeds 00, $1$, and $2$.
@@
  + Mitigation 29.55 14.09 45.94 29.86 (-20.86) (-${}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\text{\quad(-20.86)}}}$
@@
  + Mitigation 15.50 34.90 62.26 37.55 (+12.97) ${}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{\quad(+12.97)}}}$
@@
  + Mitigation 3.15 51.36 62.79 39.10 (+11.19) ${}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{\quad(+11.19)}}}$
@@
  + Mitigation 16.07 33.45 57.00 35.50 (+1.10) ${}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{\quad(+1.10)}}}$
@@
  + Mitigation 19.02 22.9 32.71 24.88 (-25.67) (-${}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\text{\quad(-25.67)}}}$
@@
  + Mitigation 122.53 16.49 99.19 79.40 (+30.04) ${}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{\quad(+30.04)}}}$
@@
  + Mitigation 69.88 47.24 83.62 66.91 (+29.38) ${}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{\quad(+29.38)}}}$
@@
  + Mitigation 70.48 28.88 71.84 57.06 (+11.25) ${}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{\quad(+11.25)}}}$
@@
  + Mitigation 34.67 58.67 83.37 58.90 (+12.79) ${}_{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\text{\quad(+12.79)}}}$
@@
  + Mitigation 10.79 9.3 22.09 14.06 (-78.88) (-${}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\text{\quad(-78.88)}}}$
@@
  + Mitigation 23.65 20.3 43.37 29.11 (-11.28) (-${}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\text{\quad(-11.28)}}}$
@@
  + Mitigation 23.04 29.42 49.61* 34.02 (-22.79) (-${}_{{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\text{\quad(-22.79)}}}$

### v5

# White Men Lead, Black Women Help? Benchmarking and Mitigating Language Agency Social Biases in LLMs

@@ Abstract
Social biases can manifest in language agency. However, very limited research has investigated such biases in Large Language Model (LLM)-generated content. In addition, previous works often rely on string-matching techniques to identify agentic and communal words within texts, falling short of accurately classifying language agency. We introduce the Language Agency Bias Evaluation (LABE) benchmark, which comprehensively evaluates biases in LLMs by analyzing agency levels attributed to different demographic groups in model generations. LABE tests for gender, racial, and intersectional language agency biases in LLMs on 3 text generation tasks: biographies, professor reviews, and reference letters. Using LABE, we unveil language agency social biases in 3 recent LLMs: ChatGPT, Llama3, and Mistral. We observe that: (1) LLM generations tend to demonstrate greater gender bias than human-written texts; (2) Models demonstrate remarkably higher levels of intersectional bias than the other bias aspects. (3) Prompt-based mitigation is unstable and frequently leads to bias exacerbation. Based on our observations, we propose Mitigation via Selective Rewrite (MSR), a novel bias mitigation strategy that leverages an agency classifier to identify and selectively revise parts of generated texts that demonstrate communal traits. Empirical results prove MSR to be more effective and reliable than prompt-based mitigation method, showing a promising research direction. We release our source code and data at https://github.com/elainew728/labe-agency.
White Men Lead, Black Women Help? Benchmarking and Mitigating Language Agency Social Biases in LLMs
Yixin Wan  and Kai-Wei Chang University of California, Los Angeles {elaine1wan,kwchang}@cs.ucla.edu
@@ 1 Introduction
Social biases manifest through drastically varying levels of agency in texts describing different demographic groups Grimm et al. (2020); Polanco-Santana et al. (2021); Stahl et al. (2022); Wan et al. (2023a). For instance, bias arises from portraying demographic minority groups—such as Black individuals and women—as being communal (e.g. “warm” and “helpful”), and dominant social groups—such as White individuals and men—as being agentic (e.g. “authoritative” and“in charge of” things) Cugno (2020); Grimm et al. (2020). However, there lacks a comprehensive benchmark for evaluating such bias in language agency. Additionally, previously proposed approaches to measure language agency are mostly limited to string matching and simple sentiment-based approaches, resulting in a lack of accuracy and reliability in agency classification results. A qualitative example is provided in Figure 1: while differences in language agency are observable in the texts, string matching yields 00 agentic and 00 communal words; a sentiment classifier labels both texts as “positive”.
@@
To address the research gaps, we propose a novel Language Agency Bias Evaluation (LABE) benchmark for comprehensively measuring gender, racial, and intersectional language agency biases in LLMs. Using 5,400 template-based entries, an accurate language agency classifier, and interpretable metrics for each bias dimension, LABE examines agency-related biases on 3 text generation tasks for LLMs: biography, professor review, and reference letter generation. For building the accurate and reliable automated agency classification tool, we also collect and contribute the Language Agency Classification (LAC) dataset with 3,724 agentic and communal sentences. Using LAC, we trained an agency classifier (achieving 91.69% test accuracy) and incorporated it into LABE to evaluate language agency biases in 3 recent LLMs: ChatGPT, Mistral, and Llama3. We observed that:
@@
Based on these observations, we further propose the Mitigation via Selective Rewrite (MSR) method for more effective and targeted mitigation of language agency biases. MSR utilizes the agency classifier to identify and revise communal sentences in model generations, yielding more agentic updated texts. Experiments show that compared to prompt-based mitigation, MSR achieves more effective and stable bias reduction results. Our LABE benchmark, LAC dataset, and the MSR mitigation method make valuable technical contributions, and introduce language agency bias as a novel direction in NLP fairness research.
@@ 2 Related Work
2.1 Language Agency in Texts
While a body of works in social science (Akos and Kretchmar, 2016; Grimm et al., 2020; Polanco-Santana et al., 2021; Park et al., 2021) and NLP (Sap et al., 2017; Ma et al., 2020; Park et al., 2021; Stahl et al., 2022; Wan et al., 2023a) studied language agency, they suffer from 2 drawbacks:
Firstly, existing works fail to establish a comprehensive evaluation benchmark for language agency biases in LLMs. Most works studied such biases in specific human-written texts (e.g. only biography), and only focused on single dimensions of bias (e.g. only gender bias), limiting the scope of analysis. As more real-world downstream applications of LLM-generated texts arise, it is critical to identify and quantify potential agency-related fairness issues in LLM generations.
Secondly, existing methods to measure language agency struggle with achieving accuracy and reliability. Prior works often utilized string matching for words in agentic and communal lexicons to measure agency. However, string matching and sentiment-based approaches only yield $46.65$ and $52.28$ in agency classification accuracy, respectively (as shown in Appendix B, Table 12).  Wan et al. (2023a) utilized a model-based agency measurement method, but only achieves $66.49\%$ classification accuracy (Appendix B, Table 12).
@@ 2.2 Biases in Human-Written and LLM-generated Texts
Bias in Reference Letters    Trix and Psenka (2003); Cugno (2020); Madera et al. (2009); Khan et al. (2021); Liu et al. (2009); Madera et al. (2019), and Wan et al. (2023a) uncovered gender biases in letters of recommendation. For instance,  Trix and Psenka (2003); Madera et al. (2009) and  Madera et al. (2019) studied bias in the “exellency” of language.  Morgan et al. (2013); Akos and Kretchmar (2016); Grimm et al. (2020); Powers et al. (2020); Polanco-Santana et al. (2021); Chapman et al. (2022); Girgis et al. (2023) investigated racial biases in reference letters:  Girgis et al. (2023) studied biases in emotional words and language traits like tone, but did not open-source their evaluation tools;  Akos and Kretchmar (2016); Grimm et al. (2020); Powers et al. (2020); Chapman et al. (2022); Polanco-Santana et al. (2021), and Chapman et al. (2022) used string matching for word-level bias analysis. For example,  Powers et al. (2020) and  Chapman et al. (2022) showed that racial minority groups are significantly less frequently described with standout words than their White colleagues.
@@ 2.3 Bias in Language Agency
An increasing body of recent studies have investigated biases in intricate language styles, such as language agency Sap et al. (2017); Ma et al. (2020); Stahl et al. (2022); Wan et al. (2023a).  Akos and Kretchmar (2016); Sap et al. (2017); Ma et al. (2020); Grimm et al. (2020); Polanco-Santana et al. (2021); Park et al. (2021), and Stahl et al. (2022) measured language agency by string matching for agentic and communal verbs, and then calculate their occurrence frequencies. However, string-matching methods fail to consider the diversity and complexity of language, and could not capture implicit indicators of language agency, as illustrated in Figure 1.  Wan et al. (2023a) was the first to adopt a model-based method to measure language agency gender biases in LLM-generated reference letters. Nevertheless, their model lacks accuracy in sentence-level classification, and the scope of their analysis is constrained to LLM-synthesized reference letters.
@@ 3 The Language Agency Bias Evaluation (LABE) Framework
Agentic language depicts “proactive” characteristics such as speaking assertively, influencing others, and initiating tasks; communal language portrays “reactive” characteristics like caring for others, providing assistance, and sustaining relationships (Madera et al., 2009; Wan et al., 2023a). We define “language agency bias” to be the unequal representation of language agency in texts depicting different demographic groups, e.g. by showing women as submissive and powerless and men as assertive and dominant (Stahl et al., 2022), or by describing racial minority groups with more communal language than agentic (Grimm et al., 2020; Polanco-Santana et al., 2021).
@@ 3.1 Generative Discriptive Texts for Demographic Groups with LLMs
Wan et al. (2023a) proposed the Context-Less Generation (CLG) setting, in which they adopt templates and descriptors to prompt for a variety of LLM-generated reference letters for different genders. Inspired by CLG, we extend the setting to 3 different text generation tasks: biography, professor review, and reference letter generation. We combine descriptors with demographic information—such as race, gender, or intersectional identities—and template-based prompts to query for LLMs’ generation. Each prompt must contain race and gender descriptors. For the name descriptor, we prompt ChatGPT to generate 5 popular names for each gender and race intersectional group. Additional descriptors like occupation and department are included to improve prompt variability. The final LABE benchmark tests LLMs on 2,400 templated-based prompts for biography generation, 600 for professor review, and 2,400 for reference letters. Note that entry numbers differ due to the difference in descriptors used (departments for professor review, whereas occupations for the other 2). Full details are in Appendix A.
@@ 3.2.1 Dataset Collection
To ensure the trustworthiness of the constructed dataset, we adopt a novel dataset construction framework that consists of an automated component and a human-involved component.
We begin by preprocessing a personal biography dataset (Lebret et al., 2016) into sentences, aiming at using these as seed texts to construct agentic and communal texts through paraphrasing. This step ensures the fairness of collected dataset, since (1) the raw data output would be balanced between the two labels, and (2) each sentence in each biography would have an agentic paraphrase and a communal paraphrase, preventing social bias propagation like having more agentic sentences for dominant social groups.
Next, we adopt Openai’s gpt-3.5-turbo-1106 model (OpenAI, 2022) to paraphrase each sentence into an agentic version and a communal version. This ensures scalability through an automated generation pipeline, and also guarantees consistency since all paraphrases would come from a single source (in contrast with using human-written paraphrases, which is hard to scale and might result in drastically subjective writing tones).
Furthermore, we utilize a human verification step to ensure the naturalness of the generated dataset. We invite $2$ human annotators, who are native speakers of English, to re-label each data and identify ambiguous cases.
Finally, data entries with ambiguity are removed and ground truth labels of the LAC dataset are decided by a majority vote between the annotators’ labels and the paraphrasing target (i.e. whether a sentence was generated as an “agentic” or “communal” paraphrase). Full details of dataset construction are in Appendix B. Details on dataset statistics are in Appendix B.5
@@ 4 Unveiling Language Agency Biases in LLMs with LABE
We utilize LABE to measure gender, racial, and intersectional biases in 3 recent LLMs. In this section, we provide details on evaluated models, observed outcomes, as well as result analysis of our evaluation experiments.
Models and Generation Settings   We experiment with 3 recent LLMs: the gpt-3.5-turbo-1106 version of OpenAI s ChatGPT (OpenAI, 2022), Llama3-8B-Instruct (Touvron et al., 2023), and Mistral-7B-Instruct-v0.2. We utilize ChatGPT’s API for experiments, with no license information. Llama3 is licensed under the Meta Llama 3 Community License and Mistral is under Apache License 2.0; both models are publicly available. For ChatGPT, we followed all default generation settings in the API call. We use Huggingface’s text generation pipeline to implement Llama3 and Mistral, and follow all default generation hyperparameters besides setting maximum number of new tokens to 512. We provide the prompts used for querying LLM generations for different tasks in Appendix E, Table 10. All results are averaged on random seeds 00, $1$, and $2$.